# TESSERA: Temporal Embeddings of Surface Spectra for Earth Representation and Analysis

### v1

@@ Abstract
Satellite remote sensing (RS) enables a wide array of downstream Earth observation (EO) applications, including climate modeling, carbon accounting, and strategies for conservation and sustainable land use. We present TESSERA, a novel Remote Sensing Foundation Model (RSFM) that uses Self-Supervised Learning (SSL) to generate global, robust representations at 10m scale from pixel-level satellite time series data. TESSERA combines information from only optical and SAR data streams using two parallel Transformer-based encoders: one dedicated to Sentinel-1 SAR polarizations and another to Sentinel-2 MSI data (10 selected spectral bands) to create representations that are then fused using a multilayer perceptron (MLP), resulting in a global representation map covering the years 2017 to 2024. Our precomputed representations set a new state-of-the-art performance benchmark and our open-source approach democratizes access to high-performance, high-resolution representations. We benchmark the performance of TESSERA in five diverse tasks, comparing our work with state-of-the-art task-specific models and other foundation models. Our results show that TESSERA outperforms both traditional RS baselines and the leading geospatial foundation models in these diverse downstream tasks.
@@
Monitoring Earth’s dynamic systems through satellite Earth Observation (EO) is critical for addressing global challenges including food security, biodiversity loss, climate change, and disaster mitigation. Petabytes of EO data are available from sensors in various modalities, and declining launch costs promise exponential growth in the future. However, practitioners face a fundamental bottleneck: the scarcity of large, accurately labeled datasets required for supervised training. Direct use of EO data also presents challenges as the data quality is affected by cloud cover, atmospheric effects, sensor biases, and non-uniform temporal sampling, preventing full utilization of this valuable resource.
@@
Users of EO data usually mitigate cloudiness by annual or seasonal compositing and then use supervised training of a neural network to map from composited images to tasks-specific classification, such as crop identification. However, compositing removes the critical temporal signal and training a task-specific neural network is computationally onerous. More recent EO approaches tackle the second problem by adopting the paradigm of foundation models, creating general-purpose models pre-trained on a broad set of EO data. For example, Visual Foundation Models (VFMs) have been engineered to interpret a wide array of geospatial data types, including multispectral, hyperspectral, and SAR imagery. Their development follows several paths, including supervised pre-training on large labeled datasets for spatiotemporal analysis [13, 12, 83, 11], and various self-supervised learning (SSL) strategies that learn from the inherent structure of the data itself. These SSL techniques include generative methods, such as masked autoencoders [19, 55, 69, 32, 62, 48, 61, 28, 77, 74, 30, 72, 37, 79] and diffusion-based models [92, 42, 64], and contrastive methods which learn efficient representations by comparing data samples [7, 54, 5, 52, 43, 14, 29, 38, 85]. Subsequently, these general-purpose models are adapted and fine-tuned for specific remote sensing tasks such as high-resolution segmentation [84, 87, 88, 16, 91].
@@
Whilst existing SSL approaches offer rich multi-dimensional ‘representations’ from unlabeled data through auxiliary tasks that reveal latent patterns [8, 51], they still do not address the problem of loss of the temporal signal during compositing. In recent work Lisaius et al [49] have demonstrated that the Barlow Twins method [90] is an elegant approach, grounded in redundancy reduction principles [10], to preserve the temporal signal in latent representations. Building on this insight, we introduce TESSERA, a foundation model for EO data that creates 128-dimensional latent representations at global scale to enable state-of-the-art performance across a diverse array of complex tasks. Our experiments demonstrate that TESSERA-learned representations outperform baseline methods in tasks as diverse as estimating canopy height in rainforests in Borneo, detecting crop types in Austria, and identifying fire scars in California. Crucially, our model is open source, enabling ease of adoption and reproducibility. Our global representation map also enables users to keep their data private rather than having to upload it to a centralized repository.
1 Our Model
TESSERA processes unlabeled time series from Sentinel-1 SAR and Sentinel-2 Multispectral Instrument (MSI) data. It is a remote sensing foundation model designed for pixel-wise feature extraction from Sentinel-1 and Sentinel-2 imagery. It comprises two main components: a dual-branch encoder and a projector network (see Fig. 1 for a detailed diagram). For each 10-meter pixel globally, it generates a compact, 128-dimensional embedding that encapsulates that pixel’s annual temporal and spectral characteristics. A primary goal of this work is the production of global, annual, 10-meter resolution representation maps covering the years 2017 to 2024. These pre-computed maps significantly lower the barrier of entry for a wide array of downstream Earth observation applications, by providing readily usable, information-rich features. The core methodology is based on a self-supervised learning paradigm, leveraging the complementary information from optical and SAR data streams.
Figure 1: Architecture and data processing pipeline of the TESSERA foundation model. a, Overview of the end-to-end workflow. Spatially aligned time-series data from Sentinel-2 MSI and Sentinel-1 SAR are first extracted for each pixel location, forming modality-specific ‘d-pixels’. For self-supervised training, two distinct augmentations are created from each d-pixel and processed through a dual-branch architecture with shared weights within each branch. The resulting 128-dimensional representations from both modalities are fused by a multilayer perceptron MLP. The fused features are then expanded to 16,384 dimensions by a large projector network. A modified Barlow Twins loss ($\mathcal{L}_{BT}+\mathcal{L}_{MIX}$) is calculated on the cross-correlation matrix of these projected features, optimizing the model to learn invariant and decorrelated representations. b, Detailed structure of a single encoder branch. A d-pixel, represented as a matrix of temporal steps versus spectral bands (e.g., 40 time-steps $\times$ 2 bands for Sentinel-1), is first embedded. A DOY-based temporal positional encoding is added to these embeddings. The resulting sequence is then processed by an 8-block Transformer Encoder. Finally, a GRU pooling layer aggregates the temporal features to produce a single 128-dimensional representation for that modality, which is then passed to the fusion MLP shown in panel a.
@@ 1 Our Model
The fundamental input unit for TESSERA is termed the “d-pixel”. For multi-spectral (from Sentinel-2) or SAR backscatter (from Sentinel-1) observations each d-pixel uniquely represents a single geographic 10-meter pixel by structuring the repeat observations from the complete annual time series of the modality into a two-dimensional array (i.e, timesteps by channels). This format explicitly preserves the temporal evolution and spectral/backscatter signatures inherent in the satellite data. The d-pixel design also intrinsically accommodates common observational gaps, such as those due to cloud cover, by masking invalid data points, which are then handled during subsequent processing stages. See Appendix A.4 and Supplementary Fig. 1.
1.1 Dual-encoder
TESSERA employs a multi-modal architecture to process and integrate information from the Sentinel-1 and Sentinel-2 observations. The model consists of specialized encoders for each modality followed by a projection head for self-supervised learning.
@@ 1.1 Dual-encoder
The model features two separate, parallel Transformer-based encoders: one for Sentinel-1 SAR backscatter data (VV and VH polarizations) and another for Sentinel-2 MSI data (10 selected spectral bands). Each encoder processes the temporal sequence of a d-pixel, utilizing multi-head self-attention mechanisms to capture complex temporal dependencies and patterns specific to each data modality. Positional encodings derived from the Day-of-Year (DOY) are incorporated to provide temporal context. An attention-pooling layer within each encoder aggregates the temporal features into a fixed-size vector representing the annual signature for that modality. See Appendix A.7 and Supplementary Figure 1 for architectural details.
1.2 Projector
@@
The modality-specific representations derived from the dual encoders are first fused using a multi-layer perceptron (MLP). This fused representation, which serves as the final 128-dimensional pixel embedding for downstream tasks, is then quantized before being fed into a high-dimensional projector network. This projector, a significantly larger MLP, expands the dimensionality of the fused representation [^1] This is a critical step inspired by Barlow Twins [90] to facilitate effective redundancy reduction during the self-supervised learning phase.
@@ 1.3 Model Training
The TESSERA model learns its representations through a self-supervised pretraining process on a large-scale, unlabeled global dataset. It is trained by optimizing a modified Barlow Twins loss function [90]. The core principle is to learn embeddings that are invariant to differing or incomplete views of the same pixel’s data (e.g., different sets of cloud-free observation dates) while minimizing redundancy between the learned feature dimensions. This is achieved by generating two distinct augmented views for each input d-pixel, primarily through independent sparse temporal sampling of annual Sentinel-1 and Sentinel-2 observations, rather than applying artificial distortions or augmentations. The model then learns to reconcile these partial views. The loss function includes the standard Barlow Twins terms for invariance and redundancy reduction, augmented with an additional mix-up regularization term to enhance robustness and inter-sample interaction during training (see Appendix A.7.2 for the full loss function, Equation 12).
The pretraining phase utilizes a vast global dataset comprising approximately 0.8 billion d-pixels, sampled from Sentinel-1 and Sentinel-2 imagery. A crucial element for successful generalization was a specific multi-level data sampling, shuffling, and batching strategy. This ensured that each training mini-batch contained a highly diverse set of d-pixels from various geographic locations and acquisition conditions, preventing model overfitting and promoting the learning of universally applicable features. The model was trained for one epoch over this dataset using distributed training techniques.
1.4 Model Inference and Global Representation Maps
Following pretraining, the TESSERA dual-encoder (with frozen weights and excluding the projector) is used to generate the final 128-dimensional representation for every 10-meter pixel globally, for each year from 2017 to 2024. This involves processing the annual Sentinel-1 and Sentinel-2 time series for each pixel through the trained encoders. The primary output of this inference stage is a set of annual, global, 10-meter resolution TESSERA representation maps. These maps are designed as readily usable, multi-channel geospatial data layers. This “Model-as-Data” approach significantly lowers the barrier to entry for end-users, as these rich, pre-computed features can be directly ingested by downstream models without the need for raw satellite data processing or running the TESSERA model itself.
@@ 2 Downstream Tasks
The effectiveness and generalizability of the TESSERA representations are rigorously evaluated across a diverse range of downstream remote sensing tasks. These include, but are not limited to, pixel-wise classification (e.g., crop type mapping), pixel-wise regression (e.g., canopy height estimation), and patch-based dense predictions (e.g., semantic segmentation and land change detection). For all such evaluations, the pretrained TESSERA encoder acts as a fixed feature extractor. Lightweight, task-specific model heads (e.g., shallow MLPs or UNet-style decoders) are then trained using these fixed representations. This approach typically requires substantially less labeled data and computational resources compared to training deep models from scratch, highlighting the transfer learning capabilities of TESSERA. See Appendix A.9 and Supplementary Fig. 9 for the general downstream application workflow.
2.1 Crop Type Classification Evaluation
To validate the effectiveness of TESSERA representations for agriculture, we performed a rigorous evaluation on a crop type classification task. We used the official Austrian INVEKOS dataset for the 2021-2022 growing season [1], which contains 17 aggregated crop classes.
Our evaluation methodology centered on comparing the performance of a lightweight classification head trained on TESSERA representations against two key baselines: (1) a traditional Random Forest classifier trained on engineered temporal features from raw satellite data, and (2) the PRESTO foundation model [79], where we applied the same lightweight head to its embeddings for a direct comparison. We assessed performance using multiple metrics, including F1 scores, under varying data availability scenarios, from 30% of labeled data down to one-shot learning. We also evaluated performance on a patch-based semantic segmentation task against several other foundation models. Further details on data processing, model architectures, and training protocols are provided in Section B.
2.2 Crop classification
@@
Agricultural monitoring through EO is a cornerstone of managing global food security, providing critical data to stakeholders in the agricultural sector [86]. Crop type classification inform vital agricultural decisions by allowing accurate estimates of crop area and yield, and facilitating low-cost monitoring of diseases and pests over vast regions [31]. However, this task is notably more complex than many other land cover classification challenges, as it requires discerning subtle spectral and temporal variations that differentiate various crop classes [31]. For years, classifiers like Random Forest (RF) applied to hand-crafted composite time-series data have served as a standard baseline [31, 63].
Recent advances have shifted towards foundation models, which, unlike bespoke machine learning models that require region-specific training, offer generalizable representations that can be applied universally [4, 41]. Here, we demonstrate that TESSERA’s pre-computed representations set a new state-of-the-art. We compare TESSERA against two critical benchmarks: the traditional RF baseline and the leading pixel-based foundation model, PRESTO (Pretrained Remote Sensing Transformer) [79], a transformer-based model utilizing multi-modal inputs. Our results show TESSERA not only substantially outperforms both, but also offers a more streamlined and efficient application workflow.
@@
To evaluate performance, we used the INVEKOS Austrian crop dataset from the 2021–2022 growing season, an extensive data set that covers 1850 km2 east of Vienna with 154 different crop types, which we consolidated into 17 classes based on phenology and data availability [1]. TESSERA consistently surpasses both baselines across all training regimes, especially in low-data settings crucial for operational deployment where labeled data is scarce (Fig. 2). In pixel-wise classification tasks with training data splits from 1% to 30%, a simple multilayer perceptron (MLP) trained on TESSERA representations achieves higher average and balanced F1 scores than both RF and an identical MLP trained on PRESTO embeddings, with performance gains frequently exceeding 10% and 30%, respectively (Fig. 2a). TESSERA also maintains statistically significant advantages in one-shot and few-shot scenarios (Fig. 2b). Furthermore, TESSERA’s utility extends to patch-based semantic segmentation, where it consistently achieves higher mIoU and macro F1 scores compared to other leading foundation models across various patch sizes (Fig. 2c).
@@
These performance gains are rooted in the superior quality of the learned representations. A UMAP analysis of the embedding space reveals that TESSERA’s representations form tighter, more semantically coherent clusters by crop class compared to those from PRESTO (Fig. 2d). This improved separability highlights TESSERA’s strength as a foundational tool for agricultural applications. Notably, using TESSERA’s compact, pre-computed 128-dimensional representations eliminates the need for the complex data preparation and feature engineering required by traditional methods like RF, democratizing access to high-performance crop classification.
@@
Figure 2: TESSERA representations achieve state-of-the-art performance in crop type classification. The evaluation was conducted on the 2022 Austrian INVEKOS dataset. a, Average and Balanced F1 scores for pixel-wise classification as a function of the training data ratio. TESSERA embeddings, coupled with a simple MLP, consistently outperform a Random Forest model trained on raw time-series data and the PRESTO foundation model. b, Overall F1 score in few-shot learning scenarios. TESSERA maintains a significant performance advantage even with very few training samples per class. Error bars represent the standard deviation over multiple runs. c, Patch-based semantic segmentation performance comparison. TESSERA representations, used with a UPerNet head, achieve higher mean Intersection over Union (mIoU) and Balanced F1 scores than other foundation models across various patch sizes and training data ratios. d, UMAP visualization of the embedding spaces for TESSERA and PRESTO for 17 crop classes. TESSERA’s embeddings exhibit clearer separation and more coherent clustering, as supported by superior Silhouette and Davies-Bouldin scores, indicating a more semantically meaningful representation space.

### v2

@@ Abstract
Satellite remote sensing from repeated observations and multiple sensors enables a wide range of downstream applications, including climate modeling, carbon accounting, and strategies for conservation and sustainable land use. However, satellite time series are voluminous, often corrupted by sensor noise, clouds, and atmospheric conditions, and unevenly spaced in time, making them challenging to use. We present TESSERA, an open, global, land-oriented remote sensing foundation model that uses self-supervised learning to generate ‘ready-to-use’ embeddings at 10 m scale from pixel-level satellite time series data. TESSERA uses two parallel Transformer-based encoders to combine optical data from ten Sentinel-2 spectral bands at 10-60 m spatial resolution and two Sentinel-1 synthetic aperture radar backscatter coefficients at 10 m resolution to create embeddings that are subsequently fused with a multilayer perceptron to create annual global embedding maps. We compare our work with state-of-the-art task-specific models and other foundation models in five diverse downstream tasks and find that TESSERA closely matches or outperforms these baselines. We believe that TESSERA’s ease of use, openness, computation-, label-, and data-efficiency, and high performance will prove transformative in a wide range of ecological applications.
@@
Monitoring Earth’s dynamic systems through satellite Earth Observation (EO) is critical to addressing global challenges including food security, forest and biodiversity loss, climate change, and disaster mitigation [118, 183, 24]. Petabytes of EO data are available from sensors in various modalities, and declining launch costs promise exponential growth in the future [176]. However, practitioners face a fundamental bottleneck: the scarcity of large, accurately labeled datasets required for supervised training [95]. The direct use of EO data also presents challenges as data quality is affected by sensor noise, cloud cover, atmospheric effects, and nonuniform temporal sampling, preventing full and easy use of this valuable resource [193, 60, 130].
@@
EO data users usually mitigate these issues by annual or seasonal compositing [34], then using supervised machine learning to map from composited images to task-specific classification, such as crop identification [8]. However, compositing removes the critical temporal signal arising from phenology and training a task-specific ML model is computationally onerous. Recent EO approaches tackle the latter problem using foundation models: general-purpose models pre-trained on a broad set of EO data [178]. Their development follows several paths, including supervised pre-training on large labeled datasets for spatio-temporal analysis [21, 20, 160, 18], and various self-supervised learning (SSL) strategies that learn from the inherent structure of the data itself. These SSL techniques include generative methods, such as masked autoencoders [32, 104, 126, 61, 113, 94, 112, 53, 144, 140, 134, 67, 149], diffusion-based models [192, 80, 115], and contrastive methods that learn efficient representations by comparing data samples [14, 102, 7, 99, 81, 23, 59, 70, 169]. Subsequently, these general purpose models are adapted and fine-tuned for specific remote sensing tasks such as high-resolution (semantic) segmentation [161, 177, 184, 28, 191].
@@
Although existing SSL approaches create rich multidimensional ‘embeddings’ from unlabeled data through auxiliary tasks that reveal latent patterns [15, 97], they still do not address the loss of the temporal signal during compositing. In recent work, Lisaius et al [95] have demonstrated that the Barlow Twins loss [187], grounded in redundancy reduction principles [17], can be adapted to preserve the spectral-temporal signal in latent embeddings when processing dense EO time series. Building on this insight, we introduce TESSERA, a pixel-oriented foundation model for EO time series that creates 128-dimensional latent embeddings at 10m resolution and global scale to enable state-of-the-art performance across a diverse array of complex tasks. Our experiments demonstrate that these TESSERA-derived embeddings outperform baseline methods in tasks as diverse as estimating canopy height in rainforests in Borneo, detecting crop types in Austria, and identifying fire scars in California.
Importantly, our outputs are ‘ready to use’: Users need neither to fine-tune the model, nor download and process voluminous satellite data, nor perform inferencing other than what is needed to perform a downstream task. This makes TESSERA frugal in terms of both data handline and computation. They can also keep their data private instead of having to upload it to a centralized repository. This removes a crucial barrier to use. We demonstrate that TESSERA is label-efficient: Only a relatively small number of labels for task-specific training are needed to achieve state-of-the-art performance. Moreover, our model is open source, enabling extensibility and scientific reproducibility.
We have made a gap-free, global, annual, 10-meter resolution embedding map for 2024 openly available through the Geotessera Python library [152] and aim to eventually provide global maps for the entire period 2017-2024. These information-rich precomputed maps greatly simplify a wide range of downstream EO applications. We have also released interactive and batch tools for habitat mapping using TESSERA embeddings (see [1] for details).
To our knowledge, TESSERA is unprecedented in its ease of use, scale, and accuracy: no other foundation model provides analysis-ready outputs, is open, and provides global, annual coverage at 10m resolution using only spectral-temporal features at pixel level (Supplementary Text:  LABEL:tab:sup_comparison).
1 Overview of the TESSERA System
TESSERA uses self-supervised learning to create per-pixel embeddings from the unlabeled time-series of observations from Sentinel-1 (SAR) and Sentinel-2 Multispectral Instruments (MSI), learning complementary information from the optical and microwave data streams. It comprises two main components: a dual-branch encoder and a projector network (see Fig. 1 for details).
@@ 1 Overview of the TESSERA System
TESSERA models its input as a set of ‘d-pixels.’ Each d-pixel represents the annual time series of Sentinel-1 and Sentinel-2 observations at a specific 10 m pixel on Earth as a two-dimensional array ordered in time (that is, channels $\times$ timesteps). This format preserves the temporal evolution of spectral/backscatter signatures inherent in the satellite data. Our design accommodates common observational gaps and irregularities, such as those due to cloud cover or orbit overlap, by masking invalid rows, which are handled during subsequent processing stages. Model training and inferencing is done by sparse temporal sampling followed by flattening of these d-pixels (see Appendix A.4 and Fig. 1 for details). For each 10 m pixel, the fused TESSERA encoders generate a compact (8 bit), 128-dimensional embedding that encapsulates that pixel’s annual temporal and spectral characteristics. Under typical mid-latitude observation conditions, this compresses input data by a factor of $\approx$ 24, with higher compression ratios at higher latitudes.
Figure 1: Architecture and data processing pipeline of the TESSERA foundation model. a, Overview of the end-to-end workflow. Spatially aligned time-series data from Sentinel-2 MSI and Sentinel-1 SAR are first extracted for each pixel location, forming modality-specific ‘d-pixels’. For self-supervised training, two distinct augmentations are created from each d-pixel–each augmentation is a random cloud-free sample from an annual time series of Sentinel-2 observations along with the corresponding Sentinel-1 data–processed through a dual-branch architecture with weights shared between the branches. The resulting 128-dimensional embeddings from both modalities are fused by a multilayer perceptron (MLP) to create the output embeddings of the system. The fused features are then expanded to 16,384 dimensions by a large projector network. A modified Barlow Twins loss ($\mathcal{L}_{BT}+\mathcal{L}_{MIX}$) is calculated on the cross-correlation matrix of these projected features, optimizing the model to learn augmentation-invariant and decorrelated embeddings that are internally decorrelated, and fairly invariant to obscured days in the Sentinel-2 time series. b, Detailed structure of a single encoder branch. A d-pixel, represented as a matrix of temporal steps versus spectral bands (40 time-steps $\times$ 2 bands for Sentinel-1), is first embedded. A Day-of-Year-based temporal encoding is added to these embeddings. The resulting sequence is then processed by an 4-block Transformer Encoder. Finally, a Gated Recurrent Unit (GRU) pooling layer processes the sequence of temporal features; its final hidden state, which summarizes the entire time series, serves as the single 128-dimensional embedding for that modality, which is then passed to the fusion MLP shown in panel a.
@@
TESSERA uses two separate, parallel Transformer-based encoders: one for Sentinel-1 SAR backscatter data in the C-band (VV and VH polarizations) and another for Sentinel-2 MSI data (10 spectral bands covering the visible, near, and shortwave infrared). Each encoder processes the temporal sequence of a d-pixel, utilizing multi-head self-attention mechanisms to capture complex temporal dependencies and patterns specific to each data modality. Positional encodings derived from the Day-of-Year (DoY) are incorporated to provide temporal context. An attention pooling layer within each encoder aggregates the temporal features into a fixed-size vector that represents the annual signature for that modality [1]. See Appendix A.7 and Figure 1 for architectural details.
@@
Modality-specific embeddings derived from dual encoders are first fused using a multilayer perceptron (MLP) which is trained alongside the encoders. This fused embedding, which serves as the 128-dimensional pixel embedding for downstream tasks, is then quantized before being fed into a high-dimensional projector network. This projector, a significantly larger MLP, expands the dimensionality of the fused representation to facilitate effective redundancy reduction during the self-supervised learning phase. A detailed breakdown of the model’s architectural parameters is provided in Appendix A.6 (Table 2).
@@ 1.1 Model Training
TESSERA learns the mapping of inputs to embeddings through a self-supervised pretraining process on a large-scale, unlabeled global dataset. It is trained by optimizing a modified Barlow Twins loss function [187, 95]. The core principle is to learn embeddings that are invariant to differing views of the same pixel’s optical-band data (e.g., different subsets of cloud-free observation dates) while minimizing redundancy between the learned feature dimensions. This is achieved by generating two distinct augmented views for each input d-pixel, through independent sparse temporal samples. This avoids creating augmentations by applying artificial (e.g. hand-crafted) distortions. The model then learns to reconcile these partial views. The loss function includes the standard Barlow Twins terms for invariance and redundancy reduction, complemented with an additional mix-up regularization term to enhance robustness and intersample interaction during training (see Appendix A.7.2 for the full loss function, Equation 11). In practice, we randomly sub-sample each d-pixel three times, to create a total of three pairs of augmentations.
The pretraining phase utilizes a vast global dataset comprising approximately 800 million d-pixels, down-sampled from Sentinel-1 and Sentinel-2 imagery (we use a downsampling ratio of 400:1). To ensure successful generalization, a multilevel data sampling, shuffling, and batching strategy was developed. This ensured that each training mini-batch contained a highly diverse set of d-pixels from various geographic locations and acquisition conditions, preventing model overfitting and promoting the learning of universally applicable features. The model was trained for one epoch on this dataset using distributed training techniques (see Appendix A.7.3 for further details on data preparation, the shuffling strategy and optimization). The key hyperparameters used during pretraining are summarized in Appendix A.7.3 (Table 3).
1.2 Model Inference and Global Embeddings Maps
Following pretraining, the TESSERA dual encoder (with frozen weights and excluding the projector) is used to generate an annual 128-dimensional embedding for every 10-meter pixel globally. This involves making inferences from a sample of length 40 from the annual Sentinel-1 and Sentinel-2 time series for each pixel through the trained encoders. To ensure that our maps are gap-free, random sampling with replacement is used to generate augmentations from d-pixels that have fewer than 40 non-cloudy observations. The primary output of this inference stage is a set of annual, global, 10-meter resolution TESSERA embedding maps. These maps are designed as readily usable multichannel geospatial data layers. This “Embeddings-as-Data” approach significantly lowers the barrier to entry for end-users, as these rich, precomputed features can be directly ingested by downstream models without the need for raw satellite data processing or running the TESSERA model itself (see Appendix A.8). We are generating global embeddings for 2024 (approx. 250 TB) and will progressively extend coverage backward year by year until 2017.
@@ 2 Downstream Tasks
The effectiveness and generalizability of the TESSERA embeddings are evaluated on a diverse range of downstream remote sensing tasks. These include, but are not limited to, pixel-wise classification (e.g., crop type mapping), pixel-wise regression (e.g., canopy height and AGB estimation) and patch-based dense predictions (for example, semantic segmentation). In these evaluations, the pre-trained TESSERA encoder acts as a fixed feature extractor; given adequate computing resources, fine-tuning of the model is also possible. Lightweight, task-specific model heads (e.g., shallow MLPs or UNet-style decoders) are then trained using these fixed embeddings. This approach typically requires substantially less labelled data and computational resources compared to training deep models from scratch, yet achieves or surpasses the state of the art. See Appendix D and Supplementary Fig. 10 for the general downstream application workflow.
We have studied how the model’s performance on a typical downstream task improves as we increase the projector dimension (responsible for more than 95% of the model’s parameters) and the amount of computing as well as carried out an ablation study to understand which model components and design decisions are mainly responsible for performance  [1].
2.1 Crop Type Classification
@@
Agricultural monitoring through EO is a cornerstone of managing global food security, providing critical data to stakeholders in the agricultural sector [175]. The classification of crop types informs vital agricultural decisions by allowing accurate estimates of crop area and production, and facilitating low-cost monitoring of diseases and pests in vast regions [60]. However, this task is considerably more complex than many other challenges in the classification of land cover, as it requires discerning subtle spectral and temporal variations that differentiate different crop classes [60]. In the past, classifiers such as Random Forest (RF) applied to hand-crafted composite time series data have established the standard [60, 114]. Recent advances have shifted towards foundation models, which, unlike bespoke machine learning models that require region-specific feature extraction and training, offer generalizable representations that can be applied universally [6, 79]. Here, we demonstrate that TESSERA’s precomputed embeddings set a new state-of-the-art.
To evaluate performance, we used the INVEKOS Austrian crop dataset [3] from the 2021–2022 growing season. This is an extensive data set that covers 1850 km2 east of Vienna with 154 different crop types, which we consolidated into 17 classes based on crop similarity (e.g., merging feed beet with sugar beets) [3]. We compared the performance of a lightweight classification head trained on TESSERA embeddings against three key baselines: (1) an RF classifier trained on engineered temporal features from raw satellite data, (2) the PRESTO foundation model [149], and (3) Google Satellite Embedding (GSE) [56]. For the foundation models, we applied the same lightweight head to their respective embeddings for a direct comparison. We evaluated performance using multiple metrics, including F1 scores, under varying data availability scenarios, from 30% of labelled data to one-shot learning. We also evaluated performance on a patch-based semantic segmentation task against several other foundation models. Details on data processing, model architectures, and training protocols are provided in Appendix E.
@@
TESSERA consistently surpasses both baselines in all training regimes, especially in settings where labeled data are scarce (Fig. 2). In pixel-wise classification tasks with training data splits from 1% to 30%, a simple MLP trained on TESSERA embeddings achieves higher weighted and macro F1 scores than an identical MLP trained on PRESTO or GSE embeddings, and substantially outperforms the RF baseline (Fig. 2a). TESSERA also maintains statistically significant advantages in one-shot and few-shot scenarios (Fig. 2b). Furthermore, TESSERA’s utility extends to patch-based semantic segmentation, where it consistently achieves higher mIoU and macro F1 scores compared to other leading foundation models, including PRESTO and GSE, across various patch sizes (Fig. 2c).
@@
These performance gains are due to the high quality of the learned embeddings. A UMAP analysis [103] of the embedding space reveals that TESSERA embeddings form tighter, more semantically coherent clusters by crop class compared to those from both PRESTO and GSE (Fig. 2d). This improved separability highlights the strength of TESSERA as a foundational tool for agricultural applications.
@@
Figure 2: TESSERA embeddings achieve state-of-the-art performance in crop type classification. The evaluation was conducted on the 2022 Austrian crop dataset. a, Weighted and Macro F1 scores for pixel-wise classification as a function of the training data ratio. TESSERA embeddings, coupled with a simple MLP, consistently outperform a Random Forest model and other foundation models like PRESTO and Google Satellite Embedding (GSE). b, Patch-based semantic segmentation performance comparison. TESSERA embeddings, used with a UNet model, achieve higher mean Intersection over Union (mIoU) and Macro F1 scores than other foundation models across various patch sizes and training data ratios. c, Weighted F1 score in few-shot learning scenarios. TESSERA maintains a significant performance advantage even with very few training samples per class. Error bars represent the standard deviation over multiple runs. d, UMAP visualization of the embedding spaces for TESSERA, PRESTO, and GSE for 17 crop classes. TESSERA’s embeddings exhibit clearer separation and more coherent clustering visually.